PROBLEM B: In the national achievement test, a school had a total mean of 59 with a standard deviation of 7. If 40 students who took the test are selected at random, what is the probability that the mean of their scores fall between 59 and 60?


1
Expert's answer
2022-04-17T15:15:11-0400

We have a normal distribution, "\u03bc=59,\u03c3=7,n=40."

Let's convert it to the standard normal distribution,

"\bar{z}=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}},\\
\bar{z}_1=\cfrac{59-59}{7/\sqrt{40}}=0,\\
\bar{z}_2=\cfrac{60-59}{7/\sqrt{40}}=0.90,\\
P(59<\bar{X}<60)=P(0<\bar{Z}<0.90)=\\
=P(\bar{Z}<0.90)-P(\bar{Z}<0)=\\
=0.8159-0.5000=0.3159\text{ (from z-table).}"